OpenStudy (anonymous):

write a equation that represents the problem, then solve the problem ; seven-fourths decreased by a number is 2 . find the number

OpenStudy (anonymous):

7/4 - x = 2 -x = 2 -7/4 -x = 8/4 -7/4 -x = 1/4 x= -1/4 check: 7/4 - (-1/4) =2 7/4 + 1/4 =2 8/4 = 2 QED.